# Joint-GCG: Unified Gradient-Based Poisoning Attacks on Retrieval-Augmented Generation Systems

### v1

@@ 5.1 Experimental Setup
A comprehensive experimental evaluation was conducted to rigorously assess the effectiveness of our Joint-GCG framework, encompassing both targeted and batch query poisoning scenarios. We evaluated Joint-GCG’s performance against state-of-the-art baselines – PoisonedRAG[11] (with GCG[24] on the generator), LIAR[13], and Phantom[15] – under diverse conditions. This thorough evaluation included experiments focusing on targeted query poisoning, ablation studies, synthetic corpora evaluations, black-box transferability, batch query poisoning, and extended attacking steps (Appendix E). We also investigated the impact of common defensive mechanisms on Joint-GCG (Section 5.3).
5.1.1 Datasets
MS MARCO[34]: The Microsoft Machine Reading Comprehension (MS MARCO) dataset is a large-scale benchmark for information retrieval and question answering. It consists of real-world queries sampled from Bing’s search logs, with passages extracted from web documents as candidate answers. Our experiments use some subsets of the queries determined by the baselines we compare.
Natural Questions (NQ)[35]: This dataset consists of naturally occurring questions posed by users in Google Search, with human-annotated answers extracted from Wikipedia articles. Unlike MS MARCO, which focuses on search engine queries, NQ emphasizes the extraction of factual knowledge from structured sources. The dataset is particularly useful for evaluating a system’s ability to retrieve and extract concise answers from a large-scale knowledge base.
HotpotQA[36]: A multi-hop question-answering dataset that requires reasoning over multiple documents to arrive at a correct answer. Unlike single-hop QA datasets, where the answer is typically found within a single passage, HotpotQA demands integrating information across different documents, making it an excellent benchmark for assessing the model’s capability to handle complex reasoning tasks.
The dataset used for comparison with PoisonedRAG and LIAR in Experiments 5.2.1, 5.2.2, 5.2.3, and 5.2.5 consists of 100 queries, each sampled by PoisonedRAG from the MS MARCO, NQ, and HotpotQA datasets.
For comparison with Phantom in Experiment 5.2.4, we used the same dataset of 25 queries for each trigger. Each query was sampled by Phantom from the MS MARCO dataset and was selected to contain a particular trigger word.
5.1.2 Retriever Models
Contriever[37]: A contrastive learning based retrieval model designed for unsupervised sentence embeddings. Contriever is trained using a contrastive loss, which encourages similar text pairs to have closer embeddings while pushing apart unrelated pairs. This approach has demonstrated strong performance in retrieval tasks, particularly in zero-shot and low-resource settings, making it a robust choice for open-domain QA scenarios. We utilize the contriever-msmarco model here.
BGE (BAAI General Embedding)[38]: A family of embedding models developed by the Beijing Academy of Artificial Intelligence (BAAI), optimized explicitly for retrieval tasks. BGE models are trained using large-scale datasets designed to produce highly efficient representations, facilitating fast and accurate retrieval in Approximate Nearest Neighbor (ANN) search. Their effectiveness in dense retrieval tasks makes them a competitive alternative to traditional retrieval methods. We utilize the BGE-base-en-v1.5 model here.
5.1.3 Generator Models
Llama3-8B[39]: A cutting-edge open-source LLM developed by Meta AI featuring 8 billion parameters. As a successor to the highly successful Llama and Llama 2 models, Llama3 is designed to offer superior reasoning, comprehension, and response generation capabilities. Its accessibility and state-of-the-art performance make it a strong candidate for evaluating adversarial robustness in RAG scenarios.
Qwen2-7B[40]: A 7-billion-parameter model developed by Alibaba Cloud as part of the Qwen series. Qwen2 is known for its strong multilingual capabilities and efficient inference, making it a competitive choice for real-world applications. Its training methodology emphasizes knowledge-rich responses, which makes it particularly interesting to evaluate how retrieval-augmented attacks influence factual generation and reasoning.
5.1.4 Metrics
@@
Generation Attack Success Rate (): The percentage of target queries for which the LLM generates the desired target output, i.e., the generated output contains the target. We use this approach to align with PoisonedRAG[11], as it shows negligible differences from human evaluation.
@@
5.1.5 Experimental Settings
In experiments comparing Joint-GCG to PoisonedRAG with GCG and LIAR (Experiments 5.2.1, 5.2.2, 5.2.3, and 5.2.5), we set the optimizable sequence length to 32 tokens and the optimization steps to 64. We employed a variant of GCG, MCG[15], to enhance the attack efficiency and utilized its default hyperparameters. We expressly set the Batch Size to 128 and TopK to 16, used ASCII-character-only tokens for attacks, and configured the optimization target to the incorrect answer. For the LIAR method, we used a 1:1 ratio for the optimizable sequence length (16 tokens each for the retriever and generator) and optimization step count (8 steps each for the retriever and generator), attacking the retriever first, followed by the generator.
In experiment 5.2.4, we employed the same prescribed in Phantom for Denial-of-Service attacks and concatenated the optimizable sequence on it. Due to the time-consuming nature of the computation, we performed only one round of experiments. We set the optimizable sequence to 128 and the optimization step counts to 32. For Phantom, we followed their settings, using a retriever-optimizable sequence length of 128 and a generator-optimizable sequence length of 8. We perform 256 steps of GCG for the 128 retriever tokens in advance for both methods to ensure a robust retrieval rate on batch queries, facilitating further optimization of the generator-optimizable tokens.
To compare Joint-GCG with PoisonedRAG and LIAR, we used the same data as PoisonedRAG’s black-box retriever approach, using their first generated fake corpus111Available at PoisonedRAG’s official GitHub Repository. of towards the corresponding dataset in Experiment 5.2.1, 5.2.2, 5.2.3, and 5.2.5. We add an optimizable sequence at the beginning of the fake corpus, initialized with "!".
For comparisons with Phantom (Experiment 5.2.4), the documents consist of three parts concatenated together as prescribed in their work, , , and , respectively. We first initialized with "?" and with "!".
@@ 5.2 Experiment Results
5.2.1 Targeted Query Poisoning: Baseline Comparison
To evaluate the effectiveness of Joint-GCG in targeted query poisoning, we compare it against state-of-the-art poisoning methods, specifically PoisonedRAG with GCG on generator (denoted as GCG below) and LIAR. We inject one poisoned document per target query and measure the effectiveness of the attack at retrieval and generation stages for the target queries using , , and .
As shown in Table I, Joint-GCG consistently outperforms GCG and LIAR regarding and across various settings. Specifically, Joint-GCG successfully maintains near-perfect () for Llama3 and Qwen2 across all datasets. Joint-GCG significantly surpasses GCG and LIAR in attacking the generation phase, particularly on NQ and HotpotQA datasets. For instance, Joint-GCG yields up to for Llama3 and for Qwen2 on HotpotQA, outperforming GCG and LIAR by several percentage points.
@@ 5.2.1 Targeted Query Poisoning: Baseline Comparison
5.2.2 Ablation Study
To analyze the contribution of each component in Joint-GCG, we conduct an ablation study, removing key modules and evaluating their impact on attack performance.
Effect of Cross-Vocabulary Projection (CVP) and Gradient Tokenization Alignment (GTA)
We conducted an ablation study to assess the impact of the CVP and GTA modules. As they are interdependent, they were removed simultaneously. In this configuration, the optimization within our framework relies solely on generator-side gradients, as direct fusion is infeasible due to inherent mismatches in gradient shapes. The baseline is thus a variant using only generator gradients within the Joint-GCG structure.
As demonstrated in Table II, removing CVP and GTA resulted in a modest but consistent decrease in ASR. While seemingly subtle, achieving additional gains over an already effective generator-only optimization strategy is inherently challenging. This 2% absolute improvement on average represents several additional successful poisonings in the test sets. It signifies a significant enhancement in attack potency, particularly in security-critical scenarios where every successful compromise is crucial.
@@ 5.2.2 Ablation Study
Effect of Retriever-Side Loss
To further investigate the contribution of the retriever component, we conducted experiments by removing the retriever-side loss () when selecting the best candidates.
As presented in Table II, the removal of led to a pronounced decrease in across all datasets and generators. Specifically, for MS MARCO, decreased from 94.00% to 91.00% for Llama3 and from 96.33% to 92.33% for Qwen2. Similar trends were observed for NQ, with the most significant impact on HotpotQA for Llama3, where decreased from 97.33% to 91.33%. These results underscore the crucial role of the retriever-side loss in guiding the optimization process toward more potent poisoned documents.
Effect of Adaptive Weighted Fusion (AWF)
As shown in Figure 4, we experiment with different fixed retrieval-generation gradient weights and find that the AWF leads to the best performance. Significantly, AWF demonstrates superior or comparable performance to all fixed ratios in terms of . This improvement highlights AWF’s adaptive approach, effectively balancing retriever and generator optimization while dynamically adjusting the retriever weight to prevent degradation as needed.
@@
| Trigger | Attack / Step | 0 | 4 | 8 | 16 | 32 |
5.2.3 Synthetic Corpus: Removing Top-k Retrieval Dependency
One practical challenge in implementing Joint-GCG in real-world scenarios is accessing the top- retrieval results to attack the generator and calculating the stability metric for AWF. This requires querying the actual retriever during an attack, which might be restricted in specific settings. To address this limitation and explore the potential for a more self-contained attack strategy, we investigated the use of synthetic corpus generated by LLM to simulate the retrieval environment and enable AWF calculation without relying on real-time top- retrieval.
Table III reveals that Joint-GCG can effectively utilize a synthetic corpus, substantially reducing the dependency on top-k retrieval results. While some performance variance is present compared to actual corpus data, the achieved using the synthetic corpus remains commendable. For instance, even with the synthetic corpus, Joint-GCG reaches 62% with Contriever and Llama3, demonstrating a solid attack capability. This outcome underscores the practicality of Joint-GCG, enabling its application in scenarios where acquiring real-time retrieval information is challenging.
5.2.4 Extending to Batch Query Poisoning
Table IV showcases the results of this experiment. Joint-GCG consistently outperforms Phantom across all tested trigger keywords and optimization steps. Joint-GCG achieves significantly higher values at earlier optimization steps and plateaus at a higher success rate. For instance, on the “xbox” trigger, Joint-GCG reaches a high of 92% at step 4, whereas Phantom plateaus at a lower 80%. These results demonstrate Joint-GCG’s strong capability in batch query poisoning, achieving more effective and faster convergence to high attack success rates than Phantom. This highlights the versatility of Joint-GCG and its potential for broader applications in RAG system manipulation beyond targeted individual queries.
5.2.5 Evaluating Poison Generalization
A crucial aspect of assessing the practical threat posed by RAG poisoning attacks is evaluating the generalization capabilities of the generated poisons. If a poison crafted for a specific, known model (a "surrogate") proves effective against other unknown models, it validates a more realistic gray-box attack scenario. In such a scenario, an attacker would perform the computationally intensive joint optimization on an open-source model they control and then deploy the resulting poison against a target system whose exact components are unknown. We investigate two key dimensions of generalization: cross-retriever transferability and cross-generator transferability.
@@ 5.2.5 Evaluating Poison Generalization
To evaluate how poisons optimized for one retriever perform against another, we conducted experiments using Joint-GCG, where the generator model was fixed, and poisons were transferred between the Contriever and BGE retriever models. The results, illustrated in Figure 5, demonstrate notable transferability. Notably, the poisons exhibited strong cross-retriever transfer. For instance, with Llama3 as the fixed generator, poisons optimized for BGE and transferred to Contriever maintained a high of 96%, while poisons optimized for Contriever and transferred to BGE achieved an of 87%. A similar pattern was observed with Qwen2 as the generator, where BGE-optimized poisons transferred to Contriever with an of 100%, and Contriever-optimized poisons transferred to BGE with an of 80%. These findings indicate that poisons crafted by Joint-GCG can effectively compromise different retriever models. The strong transferability, particularly from BGE-optimized poisons to Contriever, suggests that some learned adversarial features are robust across retriever architectures. This enhances the practical threat, as an attacker might achieve success even without exact knowledge of the deployed retriever.
Cross-Generator Transferability
We assessed the ability of poisoned documents optimized by PoisonedRAG with GCG, LIAR, and Joint-GCG to influence other generators, including open-source models (Llama3, Qwen2) and a black-box model (GPT-4o). The comparative results are presented in Figure 6.
Regarding transferability between the open-source models Llama3 and Qwen2, Joint-GCG showed consistent transfer: poisons optimized for Llama3 achieved an of 41% on Qwen2, and poisons optimized for Qwen2 achieved an of 41% on Llama3. LIAR exhibited similar transfer from Qwen2 to Llama3 (41%) but slightly less from Llama3 to Qwen2 (35%). PoisonedRAG with GCG displayed lower transferability in both directions (33% for Llama3 to Qwen2, and 29% for Qwen2 to Llama3). Significantly, this optimization-driven transferability extends even to black-box commercial LLMs like GPT-4o[41]. While the absolute ASR against GPT-4o is lower, poisons optimized for Llama3 or Qwen2 still exhibit a noticeable increase (2%) in attack success compared to non-optimized poisons, unseen in previous attack methods. These findings underscore that RAG systems face a broader and more generalized attack surface than previously understood, as attackers can enhance cross-generator poison transferability through targeted optimization on readily available models.
5.3 Potential Defensive Mechanisms
Here, we report the outcomes of evaluating two widely used defensive strategies towards Joint-GCG: Perplexity-based filtering[42] and SmoothLLM[43].
5.3.1 SmoothLLM
| Retriever | Generator | Llama3 | Qwen2 | ||
| SmoothLLM | w/o | w/ | w/o | w/ | |
| Contriver | (w/o attack) | 77% | 72% | 71% | 66% |
| 94% | 53% | 96% | 56% | ||
| BGE | (w/o attack) | 87% | 85% | 85% | 82% |
| 87% | 47% | 92% | 41% | ||
As demonstrated in Table V, while SmoothLLM reduces Joint-GCG’s , the attack remains alarmingly potent even under this defense. For Contriever-retriever systems, Joint-GCG achieves 53% on Llama3 and 56% on Qwen2 with SmoothLLM enabled.
The persistent success of the attack under defenses highlights Joint-GCG’s robustness. Even when attenuated by SmoothLLM, the attack success rates remain comparable to undefended performance of prior methods (e.g., GCG’s around 70% in Table I). This demonstrates that Joint-GCG’s joint optimization paradigm fundamentally enhances attack survivability, signifying the challenges for defensive mechanisms and necessitating new research into retrieval-aware adversarial filtering.
5.3.2 Perplexity-Based Filtering
Perplexity-based filtering involves using the perplexity score to assess the likelihood of the corpus and filtering out documents that exceed a certain threshold. This approach aims to reduce the noise introduced by less relevant or spurious information retrieved by the system. Specifically, we computed the perplexity of the MS MARCO corpus and the corresponding Joint-GCG optimized poisons on Llama3. As shown in Figure 7, the perplexity distribution of Joint-GCG optimized adversarial examples is noticeably shifted towards higher perplexity values than the actual MS MARCO corpus. This suggests that perplexity could be a potential indicator for identifying and filtering out adversarial examples.
| Settings | ||
| Joint-GCG w/ PPL Constraint | 100.00% | 73.33% |
| w/o optimize | 100.00% | 49.00% |
Table VI presents the attack success rates. Remarkably, even when optimized with a perplexity constraint, Joint-GCG maintains a significantly higher than the baseline scenario where no optimization is performed, with a 73.33% for Joint-GCG with the perplexity constraint, compared to only 54.00% without optimization. This apparent increase demonstrates that perplexity-based filtering, in its simplest form, is inadequate against Joint-GCG. Even when adversarial examples are crafted to have perplexity values within the normal range, the attack remains potent and surpasses the baseline ASR by a large margin. This highlights the need for more advanced defenses to detect adversarial examples beyond simple perplexity thresholds.
@@ 6 Conclusion
We present Joint-GCG, a framework that elevates RAG poisoning via unified retrieval-generation gradient-based optimization. By harmonizing retrieval and generation objectives through Cross-Vocabulary Projection, Gradient Tokenization Alignment, and Adaptive Weighted Fusion, Joint-GCG overcomes the disjointed nature of prior attacks. Evaluations show at most 25% and an average of 5% higher attack success rates than state-of-the-art methods across multiple retrievers and generators, achieving greater within the same optimization steps. Ablations confirm the role of each component, while synthetic corpus tests and poison generalization experiments demonstrate the broad applicability. The framework’s potency in batch poisoning further underscores its practical threat. Joint-GCG provides a robust framework for understanding and mitigating the evolving threat landscape of RAG-based applications.
7 Limitations and Future Works
7.1 Computational Overhead
The joint optimization of retrieval and generation gradients introduces additional computational complexity compared to existing methods (as detailed in Appendix F). Primarily, the CVP module involves an offline, one-time pre-computation for each retriever-generator pair (approximately 2 hours on a single NVIDIA A6000 GPU in our setup) to train an autoencoder and derive a projection matrix. This cost is amortized over the matrix’s reuse for all subsequent attacks on that pair. While the primary additional burden is a manageable, offline pre-processing step, future work could explore more efficient optimization techniques or gradient approximation methods to reduce this overhead while maintaining the effectiveness of the attack.
7.2 Cross-Domain Generalization
While we demonstrate strong performance across multiple QA datasets, the generalization of Joint-GCG to other domains (e.g., code generation[44], medical applications[45], and tool-calling agents[46]) and different types of RAG architectures requires further investigation. Future work could explore domain-specific adaptations of the framework and evaluate its effectiveness across a broader range of applications and model architectures.
8 Ethical Considerations
@@
Appendix A Cross-Vocabulary Projection (CVP) Details
A.1 Motivation & Overview

### v2

@@ 4.1 Experimental Setup
Joint-GCG’s effectiveness was rigorously assessed against state-of-the-art baselines (PoisonedRAG with GCG (zou2023universal), LIAR (tan2024glue), Phantom (chaudhari2024phantom)) across diverse conditions, including targeted and batch (see Appendix D.1) query poisoning, ablation studies, synthetic corpora, and black-box transferability.
Datasets And Models
Evaluations utilized three open-domain Question-Answering (QA) datasets: MS MARCO (nguyen2016ms), Natural Questions (NQ) (kwiatkowski2019natural), and HotpotQA (yang2018hotpotqa). The synthetic corpus (GPT-4o-mini generated) simulated retrieval for AWF calculations for the Synthetic Corpus experiment. Two dense retrieval models, Contriever (izacard2021unsupervised) and BGE (xiao2024c), along with two LLM generators, Llama3-8B (dubey2024llama) and Qwen2-7B (yang2024qwen2technicalreport), were used.
Metrics
@@
Generation Attack Success Rate (): The percentage of target queries for which the LLM generates the desired target output, i.e., the generated output contains the target. We use this approach to align with PoisonedRAG (zou2024poisonedrag), as it shows negligible differences from human evaluation.
@@
Experiments were repeated three times. In experiments comparing Joint-GCG to PoisonedRAG with GCG and LIAR, we set the length of to 32 tokens and the optimization iterations to 64. We employed a variant of GCG, MCG (chaudhari2024phantom), to enhance the attack efficiency and utilized its default hyperparameter. We expressly set the candidate numbers of each step to 128 and token positions to 16, used ASCII-character-only tokens for attacks, and configured the optimization target to the incorrect answer. For the LIAR method, we used a 1:1 ratio for the length (16 tokens each for the retriever and generator) and optimization iterations (8 steps each for the retriever and generator), attacking the retriever first, followed by the generator.
@@ 4.2 Experiment Results
Targeted Query Poisoning: Baseline Comparison
Joint-GCG consistently outperforms GCG and LIAR in targeted query poisoning, achieving higher and across diverse settings (Table 1). Joint-GCG maintains near-perfect () for Llama3 and Qwen2 across all datasets, significantly surpassing baselines in generation attacks, particularly on NQ and HotpotQA. For instance, Joint-GCG yields up to for Llama3 and for Qwen2 on HotpotQA, outperforming GCG and LIAR by several percentage points. To isolate the impact of the optimization process itself, we also calculated the ASR on the subset of queries where a simple, unoptimized poison initially failed. The results shown in parentheses in Table 1 highlight the significant gains achieved by our method. Figure 4 further illustrates Joint-GCG’s superior efficacy, achieving comparable or higher with significantly fewer optimization steps, underscoring its enhanced efficiency and effectiveness. This performance gain stems from Joint-GCG’s integrated approach, preventing retrieval degradation and ensuring consistent poisoning efficacy.
@@
Ablation Study
An ablation study analyzed the contribution of Joint-GCG’s core components (Table 2). Removing Cross-Vocabulary Projection (CVP) and Gradient Tokenization Alignment (GTA) resulted in a modest but consistent 2% average decrease in , indicating their role in enhancing attack potency by enabling proper gradient fusion. More significantly, the removal of the retriever-side loss () led to a pronounced decrease in across all datasets and generators (e.g., a 3-5% drop on MS MARCO), underscoring its crucial role in guiding optimization towards potent poisoned documents. Furthermore, Adaptive Weighted Fusion (AWF) consistently demonstrated superior or comparable performance to all fixed retrieval-generation gradient weights (Figure 5), highlighting its effectiveness in dynamically balancing retriever and generator optimization.
@@
Synthetic Corpus: Removing Top-k Retrieval Dependency
Joint-GCG’s practicality was assessed by using a synthetic corpus to simulate retrieval, eliminating dependency on real-world top- results (Table 3). While some performance variance exists compared to real corpus data, using synthetic data remains commendable (e.g., 62% for Contriever/Llama3), demonstrating a solid attack capability in restricted settings.
Evaluating Poison Generalization
We evaluated the generalization of Joint-GCG’s poisons, which is critical for assessing real-world threats.
Cross-Generator Transferability
Joint-GCG also exhibited consistent cross-generator transferability (Figure 7) with the BGE retriever, on the MS MARCO dataset. Poisons optimized for Llama3 achieved 41% on Qwen2, and vice versa. While LIAR showed similar transfer from Qwen2 to Llama3 (41%), it was slightly less from Llama3 to Qwen2 (35%). Notably, Joint-GCG’s optimization-driven transferability extended to black-box commercial LLMs like GPT-4o, showing a noticeable 2% increase in attack success compared to non-optimized poisons, a phenomenon unseen in prior methods. These findings highlight that RAG systems face a broader, more generalized attack surface, as attackers can enhance cross-generator poison transferability through targeted optimization on readily available surrogate models.
@@
As shown in Figure 6, poisons optimized by Joint-GCG demonstrated high cross-retriever transferability, achieving 96% from BGE to Contriever and 87% from Contriever to BGE. TThis strong, bidirectional performance is highly competitive with LIAR (which achieved 94% and 86% in the same respective transfer scenarios) and PoisonedRAG + GCG (97% and 84%). Unlike the PoisonedRAG + GCG baseline which failed to achieve perfect retrieval (showing 95% and 73% ), Joint-GCG attained a 100% in its direct attacks, matching LIAR. This combination of perfect direct efficacy and strong transferability underscores the robustness and practical threat of Joint-GCG’s poisons.
@@ 5 Conclusion
We present Joint-GCG, a framework that elevates RAG poisoning via unified retrieval-generation gradient-based optimization. By harmonizing retrieval and generation objectives through Cross-Vocabulary Projection, Gradient Tokenization Alignment, and Adaptive Weighted Fusion, Joint-GCG overcomes the disjointed nature of prior attacks. Our framework consistently outperformed prior art, delivering substantial gains in attack success rates—up to 25% in certain scenarios—and demonstrating superior efficiency. Ablations confirm the role of each component, while synthetic corpus tests and poison generalization experiments demonstrate the broad applicability. The framework’s potency in batch poisoning further underscores its practical threat. Joint-GCG provides a robust framework for understanding and mitigating the evolving threat landscape of RAG-based applications.
Ethical Statement
@@
Acknowledgments
This work was supported by the National Key Research and Development Program of China (No. 2024YFF0618800), the National Natural Science Foundation of China Grant No. 62402484, the Basic Research Program of ISCAS Grant No. ISCAS-JCZD-202405, the China Southern Power Grid Company Limited ZBKJXM20240173, and the Youth Innovation Promotion Association Chinese Academy of Sciences. The authors wish to thank Mr. Zhu Qiming for his valuable insights during the conceptualization of the joint-attack methods. They are also grateful to the anonymous reviewers for their constructive feedback.
Appendix A Related Works
A.1 Retrieval-Augmented Generation (RAG) Systems
Retrieval-Augmented Generation (RAG) systems represent a significant advancement in mitigating the inherent limitations of Large Language Models (LLMs), such as knowledge cutoffs and hallucinations (lewis2020retrieval; gao2024retrievalaugmentedgenerationlargelanguage). The core principle of RAG is to ground LLM responses in external, verifiable knowledge. A typical RAG architecture involves a retriever and a generator. When a query is received, the retriever first fetches relevant documents or data snippets from a large external corpus (e.g., a vector database, enterprise knowledge base, or the internet) (asai2023retrieval). These retrieved contexts are then provided to the LLM (the generator) along with the original query. The LLM synthesizes this information to produce a more accurate, timely, and contextually appropriate response (gao2024retrievalaugmentedgenerationlargelanguage). This reliance on external corpora, while beneficial for accuracy and currency, introduces new attack surfaces, particularly corpus poisoning, which is the focus of our work.
A.2 Adversarial Attacks on Large Language Models
Large Language Models, despite their impressive capabilities, are susceptible to various adversarial attacks (weidinger2021ethical). These vulnerabilities include prompt injection, where malicious instructions are embedded in the input to hijack the model’s output (greshake2023more), and training data poisoning of the base LLM itself, which can introduce subtle biases or backdoors (chakraborty2018adversarial). These general vulnerabilities underscore the necessity for robust security measures throughout all stages of LLM development and deployment.
Gradient-Based Optimization for Adversarial Text Generation
Gradient-based optimization has become a cornerstone for crafting adversarial examples against LLMs. Early methods, such as HotFlip (ebrahimi2017hotflip), utilized gradients related to input tokens to identify minimal character-level or token-level perturbations that could alter model predictions. Simple substitution techniques (li2018textbugger) and heuristic optimizations (li2020bert) have achieved only modest success, primarily with smaller models. In contrast, gradient-guided approaches have shown superior efficacy against robust transformer architectures. The Greedy Coordinate Gradient (GCG) attack (zou2023universal) extended this by optimizing a universal adversarial suffix to elicit desired (often harmful) responses. Multi Coordinate Gradient (MCG), proposed in the Phantom work (chaudhari2024phantom), further refines this by considering multiple substitutions simultaneously for efficiency. Improved versions, such as I-GCG (jia2024improved), have also been explored. These methods demonstrate the power of gradient information but are primarily designed for direct attacks on the LLM, not the complex two-stage RAG pipeline.
A.3 Data Poisoning Attacks
Classical Data Poisoning in Machine Learning
Data poisoning is a type of malicious attack in which an attacker manipulates the training data of a machine learning model to compromise its behavior during inference (goldblum2022dataset). By injecting a small quantity of carefully crafted malicious samples into the training dataset, attackers primarily seek to degrade overall model performance, cause misclassification for targeted inputs, or implant backdoors triggered by specific inputs (shumailov2021manipulating; liu2019data).
Attacks on Information Retrieval Systems
Information Retrieval (IR) systems have also been targets of manipulation. Traditional search engines faced ”spamdexing” or ”search engine poisoning,” where malicious actors used techniques like keyword stuffing, hidden text, and link farms to artificially boost the ranking of certain web pages (gyongyi2005web; bevendorff2024google). Modern retrieval components, especially dense retrievers used in RAG, can be misled by imperceptible perturbations to documents, leading to ranking manipulation (farooq2019survey; mallen2022not). Malicious attacks can force rankers to incorrectly order documents or retrieve irrelevant content, even with minimal modifications to the corpus.
Corpus Poisoning in RAG Systems
The reliance of RAG systems on external corpora makes them uniquely vulnerable to corpus poisoning. Existing research has explored various strategies: PoisonedRAG (zou2024poisonedrag) focused on optimizing poisoned documents to maximize their retrieval probability. HijackRAG (zhang2024hijackrag) applied similar principles to prompt leaking and spam generation. However, optimizing solely for retrieval can compromise the linguistic qualities needed to effectively steer the generator, often necessitating a larger number of injected documents. Sequential optimization approaches, such as applying HotFlip for retrieval and then GCG for generation, can be suboptimal because modifications made in one stage may negatively impact the other. LIAR (tan2024glue) attempted to achieve better integration through an iterative loop but still lacks proper joint optimization, as it pre-assigns fixed optimizable lengths and optimization steps. Phantom (chaudhari2024phantom) introduced trigger-based batch poisoning, but optimized retrieval and generation sequentially, which sometimes led to retrieval failures. TrojanRAG (cheng2024trojanrag) targets the retriever model itself, a different threat model requiring greater access. In contrast, our Joint-GCG framework addresses these limitations by performing a truly joint optimization across both the retriever and the generator, marking a departure from these disjointed or sequential strategies.
Appendix B Detailed Methodology
This section expands on the technical details of the Joint-GCG framework, providing deeper insights into its core innovations.
B.1 Cross-Vocabulary Projection (CVP) Details
Motivation & Overview